\begin Let $t$ be an integer and suppose $G$ is a graph on $n$ vertices. Given an
 ordered set of $t$ vertices in $G$ and a set of $t$ distances, we find a sharp
degree condition for $G$ to contain a hamiltonian cycle with the $t$ vertices in
order with approximately the corresponding prescribed distances between them. We
 also show similar results for partitioning the graph into paths between chosen
    vertices or paritioning into cycles containing chosen vertices. \end \ \
                \begin{center} Advisor: Ron Gould\\ \end{center}